jagr2713 (jagr2713):

sqrt{2x}-sqrt{3}=sqrt{5}

jagr2713 (jagr2713):

@sleepyjess

TheSmartOne (thesmartone):

\(\sqrt{2x}-\sqrt{3}=\sqrt{5}\)

TheSmartOne (thesmartone):

There we go :D

jagr2713 (jagr2713):

corrrect

TheSmartOne (thesmartone):

Let @sleepyjess take this question. Get too 99 already :P

jagr2713 (jagr2713):

:D

OpenStudy (sleepyjess):

Add sqrt 3 to both sides first

TheSmartOne (thesmartone):

^ That is the correct direct. Ignore what I said :P

TheSmartOne (thesmartone):

direction*

TheSmartOne (thesmartone):

So @jagr2713 what do you get after adding \(\sqrt{3}\) on both sides? :P

jagr2713 (jagr2713):

Divide both sides by \[\sqrt{2x}\] ?

OpenStudy (sleepyjess):

Yes!

jagr2713 (jagr2713):

oh i think i got it, is the answer \[x=\frac{ \sqrt{3}+\sqrt{5} }{ 2 }\]

jagr2713 (jagr2713):

@sleepyjess

OpenStudy (sleepyjess):

Ye :)

jagr2713 (jagr2713):

Thanks :D